determine the equation of the line described. Put answer in the slope-intercept form, if possible.
Through (6,-1), parallel to 8x-9y=111

Answer by bfdccheerleader16(4)

first you put your line in slope intercept form which would end up to be y=8/9x-37/3 and then the next line that goes through the points (6,-1) would also have a slope of 8/9 because the lines are parallel. next you would use the points and the slope of the 2nd line to determine the y-intercept. So you would use point-slope form and get y+1=8/9(x-6) and you would get y+1=8/9x-16/3. then subtract the one and you get y=8/9x-13/3. and that is your line that goes through the points (6,-1).